OpenStudy (anonymous):

The sum of all even numbers upto 100 is---------?

OpenStudy (anonymous):

The even numbers have a common difference of 2. Your first term is 2 (I'm assuming you want to sum from numbers greater than or equal to 1 up to 100). From the formula for the nth term in an arithmetic sequence, we have, a[n] = a[1]+(n-1)d, and so, 100 = 2 + (n-1)*2 which yields n =50 - that is, you'll have 50 terms to sum. The sum of the first 50 terms is then given by, S[n] = (n/2)(a[1]+a[n]) = (50/2)(2+100) = 2550

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

No worries.